The purity of a sample of sodium oxalate, Na2C2O4, is determined by titrating with a standard solution of KMnO4. If a 0.5116-g sample requires 35.62 mL of 0.0400 M KMnO4 to reach the titration’s end point, what is the %w/w Na2C2O4 in the sample.


1
Expert's answer
2022-01-17T16:20:02-0500

Question:


The purity of a sample of sodium oxalate, Na2C2O4, is determined by titrating with a standard solution of KMnO4. If a 0.5116-g sample requires 35.62 mL of 0.0400 M KMnO4 to reach the titration’s end point, what is the %w/w Na2C2O4 in the sample.


Solution:

The reaction of the titration is:

16 H+(aq) + 2 MnO4-(aq) + 5 C2O42-(aq)  →  2 Mn2+(aq) + 8 H2O(l) + 10 CO2(g)

One can see that 5 moles of oxalate anion react with 2 moles of permanganate:

"n(C_2O_4^{2-}) = \frac{5}{2}n(MnO_4^-)" .

The number of the moles of the permanganate used to reach the titration's end point is:

"n(MnO_4^-) = cV = 0.0400\u00b70.03562 = 0.00142" mol.

Therefore, the number of the moles of oxalate in the sample is:

"n(C_2O_4^{2-}) = \frac{5}{2}·0.00142 = 0.003562" mol.

The molar mass of Na2C2O4 is 134 g/mol. Therefore, the mass of sodium oxalate in the sample was:

"m(Na_2C_2O_4) = nM = 0.003562\u00b7134 = 0.477" g

The mass percentage of the sodium oxalate in the sample is:

"w = \frac{0.477}{0.5116}·100\% =93.3\%" .


Answer: the mass percentage of the sodium oxalate in the sample is 93.3%.